The frequency of note emitted by source changes by 20%as it approaches observer.As it recedes away from him,the apparent frequency will be?

Grade:12

1 Answers

Arun
5 years ago
Dear Prema
 
n(120/100) = (v / v – vs) *n
 
1.2 V – 1.2 Vs = V
 
0.2 V = 1.2 Vs
 
Vs = V/6
 
when it recedes away
 
(100 – x / 100) * n = (V / V + Vs)*n
 
1 – x /100 = 6/7
 
x = 14.3 %
 
hence 14.3% different from actual frequency